OpenStudy (anonymous):

Triangle DEF is a right triangle. If FE = 24 and DE = 7, find DF. A. 62 B. 25 C. 625 D. 168

OpenStudy (anonymous):



OpenStudy (anonymous):

By Pythagoras Theorem: x^2 = 7^2 + 24^2 x = ?

OpenStudy (anonymous):

x= 49 + 576 x= 625